# SimBench: Benchmarking the Ability of Large Language Models to Simulate Human Behaviors

### v3

@@ Abstract
Large language model (LLM) simulations of human behavior have the potential to revolutionize the social and behavioral sciences, if and only if they faithfully reflect real human behaviors. Current evaluations are fragmented, based on bespoke tasks and metrics, creating a patchwork of incomparable results. To address this, we introduce SimBench, the first large-scale, standardized benchmark for a robust, reproducible science of LLM simulation. By unifying 20 diverse datasets covering tasks from moral decision-making to economic choice across a large global participant pool, SimBench provides the necessary foundation to ask fundamental questions about when, how, and why LLM simulations succeed or fail. We show that, while even the best LLMs today have limited simulation ability (score: 40.80/100), performance scales log-linearly with model size. Simulation performance is not improved by increased inference-time compute. We demonstrate an alignment-simulation trade-off: instruction-tuning improves performance on low-entropy (consensus) questions but degrades it on high-entropy (diverse) ones. Models particularly struggle when simulating specific demographic groups. Finally, we demonstrate that simulation ability correlates most strongly with deep, knowledge-intensive reasoning (MMLU-Pro, $r=0.939$). By making progress measurable, we aim to accelerate the development of more faithful LLM simulators.
@@ 1 Introduction
Large-scale human experiments and surveys have long been essential tools for informing public policy, commercial decisions, and academic research. Running experiments and surveys, however, is costly and time-consuming. Large language models (LLMs) can potentially address this challenge by simulating human behaviors quickly and at low cost, to complement or even substitute human studies. This prospect, alongside encouraging early evidence on the efficacy of LLMs as simulators (Aher et al., 2023; Argyle et al., 2023; Horton, 2023), has motivated a large body of recent work across many disciplines investigating the ability of LLMs to simulate human behaviors and tendencies (Binz et al., 2025; Bisbee et al., 2024; Dominguez-Olmedo et al., 2024; Manning et al., 2024; Hu & Collier, 2025; Hu et al., 2025a, inter alia).
@@
However, this rapid exploration has produced a fragmented body of evidence (Ma et al., 2024). Most studies evaluate a narrow set of LLMs on a specific task, yielding varied and sometimes contradictory results that make it difficult to draw broader conclusions (§5). The field lacks a unified framework to determine when, how, and why LLM simulations succeed, or how to train better simulators.
@@
To confront this challenge and move LLM simulation from a collection of isolated studies to a robust, reproducible science, we introduce SimBench: the first large-scale, standardized benchmark for human behavior simulation. By harmonizing 20 diverse datasets containing a single-turn, multiple-choice structure – spanning moral dilemmas, economic games, and psychological assessments across a vast global participant pool – SimBench provides an essential instrument to rigorously measure and compare simulation fidelity across models, tasks, and populations.
@@
Using this benchmark, we move beyond isolated experiments to build a comprehensive picture of LLM simulation. We structure our investigation through six research questions. We first establish a baseline, asking how well current LLMs perform (RQ1) and how characteristics like model size and inference-time compute affect their ability (RQ2). We find that even top models struggle (top score: 40.8/100), though performance scales log-linearly with size and, surprisingly, is not improved by scaling inference-time compute. Next, we explore the sources of this variance, asking how task selection (RQ3) and human response plurality (RQ4) affect fidelity. Fidelity varies substantially by task, and we establish a key alignment-simulation tradeoff: instruction tuning’s mode-seeking objective systematically improves performance on low-entropy (consensus) questions but actively harms performance on high-entropy (pluralistic) ones. A causal analysis confirms this is the net result of a beneficial instruction-following effect versus a harmful entropy-reduction effect. Finally, we investigate practical implications, asking if LLMs are better at simulating some demographic groups than others (RQ5) and to what extent simulation ability correlates with other model capabilities (RQ6). We show models struggle most with religious/ideological groups and that simulation ability correlates most strongly with deep, knowledge-intensive reasoning (MMLU-Pro, $r=0.939$).
@@
Progress in AI is only possible through rigorous evaluation, and large-scale benchmarks such as MMLU (Hendrycks et al., 2021) have significantly contributed to improvements in LLM capabilities. In the same spirit, SimBench provides the foundational infrastructure needed to move LLM simulation from a collection of ad-hoc studies to a measurable and systematic science. All of SimBench is available on GitHub and HuggingFace.
@@ 2.1 Data Curation
These criteria are complemented by a curation strategy that balances competing objectives. We prioritize novelty, favoring datasets not previously used in LLM simulation evaluation, but also ensure backward comparability by including well-established benchmarks (e.g., OpinionQA, ChaosNLI). Furthermore, we prioritize datasets with rich sociodemographic data to enable fine-grained analysis of specific subpopulations (§2.4). However, we make targeted exceptions for datasets like Jester and Choices13k, which, despite lacking demographic data, provide unique and essential task diversity.
This principled selection process yields the 20 datasets that comprise SimBench, which we list in Appendix L, providing details on participants and example questions. To demonstrate the rigor of our curation process and to serve as a resource for the community, we provide a list of datasets that were considered but ultimately excluded in Appendix B. Crucially, SimBench is fully modular by design, so that future work can easily add more datasets using the processing pipeline described in §2.3 below.
@@ 2.2 Benchmark Properties
The principled curation process results in a benchmark defined by two key properties: task diversity and participant diversity.
1) Task Diversity: SimBench covers a wide range of different tasks regarding the human behavior they measure. SimBench includes decision-making questions (e.g., in Choices13k, MoralMachine), where participants are presented with a set of actions that concern themselves, and they have to select the action they would hypothetically take. SimBench also includes self-assessment questions (e.g., in OpinionQA, OSPsychBig5), where participants are presented with a set of descriptions or attributes, and they have to select the one that best describes themselves. Further, SimBench includes judgment questions (e.g., in ChaosNLI and Jester) where participants are presented with some external object and a choice of labels, and they have to select the label they think fits best. Lastly, SimBench includes problem-solving questions (e.g., in WisdomOfCrowds and OSPsychMGKT), where participants are presented with a set of answers to a factual question, and they have to select the answer they think is correct. Consequently, LLMs have to accurately simulate several distinct types of human behavior in order to perform well on SimBench.
2) Participant Diversity: SimBench captures a rich demographic landscape spanning more than 130 different countries across six continents. While five datasets include US-based crowdworkers, the international scope of SimBench is substantial: 3 datasets (e.g., LatinoBarometro, AfroBarometer) exclusively feature participants from regions outside the US, 4 datasets (e.g., GlobalOpinionQA, TISP) draw from multi-country samples across different continents, and 2 datasets collect responses from a global pool of internet users. Importantly, 8 out of the 20 datasets employ representative sampling techniques, enhancing the ecological validity of these constituent components. To perform well on SimBench, LLMs must therefore demonstrate the ability to accurately simulate the behavior of human participants across diverse cultural, linguistic, and socioeconomic backgrounds. [^1] Note that, while some constituent datasets recruit representative samples, SimBench as a whole is not fully representative of any single population. We discuss this limitation in Appendix A.
@@ 2.3 Unifying SimBench Dataset Formats
A core contribution of SimBench is the harmonization of 20 heterogeneous datasets into a standardized format. This process ensures that LLM performance can be compared rigorously across diverse tasks and populations.
Question Normalization: We standardize all items into a multiple-choice format. For datasets with continuous scales (e.g., Likert scales), we map responses to discrete bins. We further ensure consistency by collapsing answer options where appropriate, limiting the maximum to 26 choices (though typically fewer than 10), and using the official English-language versions of all questions. [^2] We note that simulation ability may plausibly be correlated with prompt language, and encourage future work in this direction.
Response Aggregation: To evaluate group-level simulation, we standardize all data into group-level probability distributions. For the majority of our datasets, which provide raw individual-level responses, we create these distributions by aggregating the data ourselves. Post-stratification weights are applied whenever applicable (e.g. ESS). For the few datasets that are already provided in an aggregated format (e.g., GlobalOpinionQA), we process and normalize their existing statistics to conform to our benchmark’s schema.
We create the simulation targets in two ways: 1) Default Grouping. For every question in a dataset, we create a baseline target by aggregating responses from all participants. This represents the “default” population for that dataset (e.g., “US-based Amazon Mechanical Turk workers”) and is used to measure general simulation ability. 2) Specific Grouping: For datasets with rich sociodemographic data, we create more fine-grained targets by aggregating responses from participants sharing a specific attribute (e.g., age or gender). These targets are essential for evaluating a model’s ability to simulate narrower, more specific demographic groups. The available grouping variables for each dataset are detailed in Appendix L.
Each simulation target is paired with a prompt that describes the corresponding group. This entire harmonization process yields 10,930,271 unique question-group simulation targets. From this comprehensive set, we curate our final benchmark splits (§2.4) to enable robust evaluation of LLM simulation capabilities.
@@ 2.4 SimBench Splits
The full set of over 10 million simulation targets is too vast for practical evaluation. We therefore curate two distinct benchmark splits, each designed to probe a different facet of an LLM’s simulation capabilities.
@@
2) The SimBenchGrouped split contains only the five large-scale survey datasets in SimBench (AfroBarometer, ESS, ISSP, LatinoBarometro, and OpinionQA) because for these datasets we have enough participants to obtain meaningful group sizes even when selecting on a specific group attribute (e.g., age = 30-49). For each dataset, we select questions that exhibit significant variation across demographic groups, ensuring that the benchmark captures meaningful demographic differences in responses. This results in 6,343 test cases overall. For more details on the sampling process, see Appendix C. Conceptually, SimBenchGrouped measures the ability of LLMs to simulate responses from narrower participant groups based on specified group characteristics. [^3] Ideally, we would also like to measure LLM simulation ability for intersectional groups that combine multiple characteristics (e.g., female + age 30-49). However, selecting on multiple characteristics substantially decreases group size, thus increasing sampling noise in the response distributions. Reliable evaluation of intersectional group simulation ability would require datasets with more participants than we have access to.
@@ 3 Experimental Setup
Tested Models: To demonstrate the utility of SimBench and answer our six research questions (§1), we evaluate 45 state-of-the-art LLMs on SimBench. This includes both commercial and open-weight, base and instruction-tuned models [^4] We use the term “instruction tuning” to refer to any alignment related post-training, including by not limited to the narrowly defined instruction tuning, with model sizes ranging from 0.5B to 405B parameters. Table 1 shows the full list of models.
Model Elicitation: For each model, we collect predictions for the two main splits of SimBench2.4). To obtain model response distributions, we use one of two methods, depending on model type: 1) For base models, we directly extract token probabilities for each response option based on first-token logits. This is a natural way of eliciting a distribution out of an LLM, especially a base LLM. 2) For instruction-tuned models, we follow recent literature on LLM calibration and distribution prediction (Tian et al., 2023; Meister et al., 2025) and use verbalized distributions, e.g., “Option A: 30%, Option B: 70%”, elicited through prompting. We empirically validate this methodological choice in Appendix E, which provides strong evidence that verbalized distributions substantially and consistently outperform direct token probabilities for instruction-tuned models. This ensures each model class is evaluated under its optimal conditions. For implementation details and prompt formats, see Appendix D.
Evaluation Metric: To measure LLM simulation ability, we derive the SimBench score $S$ from Total Variation Distance TVD, defined as:
$S(P,Q)=100\left(1-\frac{TVD(P,Q)}{TVD(P,U)}\right)=100\left(1-\frac{\sum_{i}|P_{i}-Q_{i}|}{\sum_{i}|P_{i}-U_{i}|}\right)$ (1)
where $P$ is the human ground truth distribution, $Q$ is the LLM predicted distribution, and $U$ is a uniform distribution over all response options.
Conceptually, $S$ therefore measures how much more accurate the predictions from an LLM are than predictions from a uniform baseline model, which assigns equal probability to all response options for a given question. In other words, $S$ quantifies the advantage of an LLM simulation over the simplest possible guess. An $S$ score of 100 indicates perfect alignment between the LLM and the human ground truth distribution, while a score $\leq$0 indicates performance at or below the performance of a uniform baseline. We chose TVD as the basis for $S$ due to its symmetry, boundedness, and robustness to zero probabilities. For a comparison to alternative metrics, see Appendix F.
4 Results
4.1 RQ1: General Simulation Ability of LLMs
@@
To evaluate the general simulation ability of LLMs, we measure their overall SimBench score $S$ averaged across the two main splits of SimBench (Table 1 and Appendix Table 7).
We find that even leading LLMs struggle to simulate group-level human behaviors with high accuracy, as measured across the 20 datasets in SimBench. Claude-3.7-Sonnet is the best-performing model overall, but only achieves a score of 40.80 out of a maximum of 100 on SimBench. This score indicates that the response distributions predicted by Claude-3.7-Sonnet are, on average, closer to a uniform response distribution than to the true human response distribution. The top open-weight model, DeepSeek-R1, scores 34.52. The majority of the 45 models we test perform substantially worse still, scoring less than 20. Notably, nine models we test score below 0, indicating that their predicted response distributions are, on average, even further away from the true human response distribution than a uniform response distribution. A statistical analysis, detailed in Appendix H, confirms that the performance differences between most top-ranked models as well as within each model family are statistically significant. Overall, these results consolidate the mixed findings from prior work into a clearer, if somewhat sobering, picture. When evaluated across a diverse range of tasks and populations, today’s LLMs are still far from being consistently reliable, general-purpose simulators. The stark performance differences between models also caution strongly against the use of smaller, less capable models for simulation, many of which perform worse than a simple uniform baseline.

### v4

@@ Abstract
Large language model (LLM) simulations of human behavior have the potential to revolutionize the social and behavioral sciences, if and only if they faithfully reflect real human behaviors. Current evaluations of simulation fidelity are fragmented, based on bespoke tasks and metrics, creating a patchwork of incomparable results. To address this, we introduce SimBench, the first large-scale, standardized benchmark for a robust, reproducible science of LLM simulation. By unifying 20 diverse datasets covering tasks from moral decision-making to economic choice across a large global participant pool, SimBench provides the necessary foundation to ask fundamental questions about when, how, and why LLM simulations succeed or fail. We show that the best LLMs today achieve meaningful but modest simulation fidelity (score: 40.80/100), with performance scaling log-linearly with model size but not with increased inference-time compute. We discover an alignment-simulation tradeoff: instruction tuning improves performance on low-entropy (consensus) questions but degrades it on high-entropy (diverse) ones. Models particularly struggle when simulating specific demographic groups. Finally, we demonstrate that simulation ability correlates most strongly with knowledge-intensive reasoning (MMLU-Pro, $r=0.939$). By making progress measurable, we aim to accelerate the development of more faithful LLM simulators.
@@ 1 Introduction
Large-scale human experiments and surveys have long been essential tools for informing public policy, commercial decisions, and academic research. Running experiments and surveys, however, is costly and time-consuming. Large language models (LLMs) may help address this challenge by simulating human behaviors quickly and at low cost, to complement or even substitute human studies. This prospect, alongside encouraging early evidence on the efficacy of LLMs as simulators (Aher et al., 2023; Argyle et al., 2023; Horton, 2023), has motivated a large body of recent work across many disciplines investigating the ability of LLMs to simulate human behaviors (Bisbee et al., 2024; Dominguez-Olmedo et al., 2024; Manning et al., 2024; Binz et al., 2025; Hu & Collier, 2025, inter alia).
@@
However, this rapid exploration has produced a fragmented body of evidence. Most studies evaluate a narrow set of LLMs on a specific task, yielding varied and sometimes contradictory results that make it difficult to draw broader conclusions (§5). The field lacks a unified framework to determine when, how, and why LLM simulations succeed, or how to train better simulators.
@@
To confront this challenge, we introduce SimBench: the first large-scale, standardized benchmark for group-level human behavior simulation. By harmonizing 20 diverse datasets, which span moral dilemmas, economic games, and psychological assessments across a vast global participant pool, SimBench enables rigorous measurement and comparison of simulation fidelity across models, tasks, and populations.
@@
Using SimBench, we move beyond isolated experiments to build a comprehensive picture of LLM simulation. We structure our investigation around six research questions. We first establish a baseline, asking how well current LLMs perform at social simulation (RQ1) and how characteristics like model size and inference-time compute affect their simulation ability (RQ2). We find that top models achieve meaningful but modest simulation fidelity (top score: 40.80/100), though performance scales log-linearly with size and, surprisingly, is not improved by scaling inference-time compute. Next, we explore the sources of this variance, asking how task selection (RQ3) and human response plurality (RQ4) affect simulation fidelity. We find that fidelity varies substantially by task, and uncover an alignment-simulation tradeoff: instruction tuning’s mode-seeking objective systematically improves performance on low-entropy (consensus) questions but harms performance on high-entropy (pluralistic) questions. A causal analysis confirms this tradeoff stems from two opposing forces: a beneficial instruction-following effect and a harmful entropy-reduction effect. Finally, we ask if LLMs are better at simulating some demographic groups than others (RQ5) and to what extent simulation ability correlates with other model capabilities (RQ6). We show models struggle most with religious/ideological groups and that simulation ability correlates most strongly with deep, knowledge-intensive reasoning (MMLU-Pro, $r=0.939$).
@@
Progress in AI is only possible through rigorous evaluation, and large-scale benchmarks such as MMLU (Hendrycks et al., 2021) have significantly contributed to improvements in LLM capabilities. In the same spirit, SimBench provides the foundational infrastructure needed to move LLM simulation from a collection of ad-hoc studies to a measurable and systematic science. We note that while predicting group-level response distributions is the instrument used here for benchmarking, we view this as a foundational proxy for the broader capability of human behavior simulation. We acknowledge that full human simulation inherently includes interactive and complex dynamics not fully captured by static response distributions (see Appendix A). We have a Project Website, and all of SimBench is available on GitHub and HuggingFace.
@@ 2.1 Data Curation
We complement these criteria with a curation strategy that balances competing objectives. We prioritize novelty, favoring datasets not previously used in LLM simulation evaluation, but also ensure backward comparability by including well-established benchmarks (e.g., OpinionQA, ChaosNLI). Furthermore, we prioritize datasets with rich sociodemographic data to enable fine-grained analysis of specific subpopulations (§2.4). However, we make targeted exceptions for datasets like Jester and Choices13k, which, despite lacking demographic data, provide unique and essential task diversity.
This principled curation process yields the 20 datasets that comprise SimBench, which we list in Appendix L, providing details on participants and example questions. To demonstrate the rigor of our process, and as a resource for the community, we list datasets that were considered but ultimately excluded in Appendix B. Crucially, SimBench is fully modular by design, so that future work can easily add more datasets using the processing pipeline described in §2.3 below.
@@ 2.2 Benchmark Properties
As a whole, SimBench is defined by two key properties: task diversity and participant diversity.
1) Task Diversity: SimBench tasks are highly diverse in terms of the human behavior they capture. SimBench includes decision-making questions (e.g., in Choices13k, MoralMachine), where participants are presented with a set of actions that concern themselves, and they have to select the action they would hypothetically take. SimBench also includes self-assessment questions (e.g., in OpinionQA, OSPsychBig5), where participants are presented with a set of descriptions or attributes, and they have to select the one that best describes themselves. Further, SimBench includes judgment questions (e.g., in ChaosNLI and Jester) where participants are presented with some external object and a choice of labels, and they have to select the label they think fits best. Lastly, SimBench includes problem-solving questions (e.g., in WisdomOfCrowds and OSPsychMGKT), where participants are presented with a set of answers to a factual question, and they have to select the answer they think is correct. Consequently, LLMs have to accurately simulate several distinct types of human behavior in order to perform well on SimBench.
2) Participant Diversity: SimBench captures a rich demographic landscape spanning more than 130 different countries across six continents (see Appendix O for a full country-level breakdown). SimBench prioritizes international representation: samples from the Anglosphere West constitute only 27.9% of the data. [^1] We define the Anglosphere West as the U.S., Canada, U.K., Australia, and New Zealand. Even using a broader definition of “the West” that includes Western Europe, these nations account for less than half (45.9%) of the benchmark. This substantial global scope is driven by a diverse collection of sources: 3 datasets (e.g., Latinobarómetro, Afrobarometer) exclusively feature participants from regions outside the US, 4 datasets (e.g., GlobalOpinionQA, TISP) draw from multi-country samples across different continents, and 2 datasets collect responses from a global pool of internet users. Importantly, 8 out of the 20 datasets employ representative sampling techniques, enhancing the ecological validity of these constituent components. To perform well on SimBench, LLMs must therefore demonstrate the ability to accurately simulate the behavior of human participants across diverse cultural, linguistic, and socioeconomic backgrounds. [^2] Note that, while some constituent datasets recruit representative samples, SimBench as a whole is not fully representative of any single population. We discuss this limitation in Appendix A.
@@ 2.3 Unifying SimBench Dataset Formats
A core contribution of SimBench is the harmonization of 20 heterogeneous datasets into one standardized format. This process ensures that LLM simulation ability can be compared rigorously across diverse tasks and populations.
Question Normalization: We standardize all items into a multiple-choice format, making minimal edits that primarily consist of mapping existing discrete options to standardized letter keys, each corresponding to a single token, to enable clean extraction of per-option probabilities from base models while preserving the original experimental structure. For the few datasets with continuous scales (Jester), we map responses to discrete bins. We further ensure consistency by collapsing answer options where appropriate, limiting the maximum to 26 choices (though typically fewer than 10), and using the official English-language versions of all questions. [^3] We note that simulation ability may plausibly be correlated with prompt language, and encourage future work in this direction. This is a deliberate choice to standardize the evaluation and avoid confounding simulation ability with multilingual performance capabilities, ensuring that differences in scores reflect simulation fidelity rather than translation quality, even for datasets originally collected in local languages.
Response Aggregation: To evaluate group-level simulation, we standardize all data into group-level probability distributions. For the majority of our datasets, which provide raw individual-level responses, we create these distributions by aggregating the data ourselves. Post-stratification weights are applied whenever applicable (e.g., ESS). For the few datasets that are already provided in an aggregated format (e.g., GlobalOpinionQA), we process and normalize their existing statistics to conform to the SimBench schema.
We create the simulation targets — the ground-truth response distributions that models must predict — in two ways: 1) Default Grouping. For every question in a dataset, we create a baseline target by aggregating responses from all participants. This represents the “default” population for that dataset (e.g., “US-based Amazon Mechanical Turk workers”) and is used to measure general simulation ability. 2) Specific Grouping. For datasets with rich sociodemographic data, we create more fine-grained targets by aggregating responses from participants sharing a specific attribute (e.g., age or gender). This allows us to evaluate LLM ability to simulate narrower, more specific demographic groups. We detail the available grouping variables for each dataset in Appendix L.
Each simulation target is paired with a prompt that describes the corresponding group. Overall, this harmonization process yields 10,930,271 unique question-group simulation targets. From this set, we curate our final benchmark splits (§2.4) to enable robust evaluation of LLM simulation capabilities. We note that while training data contamination remains an inherent risk, our zero-shot aggregation task largely mitigates this by testing distributional prediction (see Appendix A for a full discussion).
@@ 2.4 SimBench Splits
The full set of over 10 million simulation targets is too vast for practical evaluation. We therefore curate two benchmark splits, each designed to probe a different facet of LLM simulation ability.
@@
2) The SimBenchGrouped split draws from the five large-scale survey datasets in SimBench (Afrobarometer, ESS, ISSP, Latinobarómetro, and OpinionQA), which are the only datasets with sufficient participant counts to yield reliable response distributions even when conditioning on a specific demographic attribute (e.g., age = 30-49). For each dataset, we select questions that exhibit significant variation across demographic groups, ensuring that the benchmark captures meaningful demographic differences in responses. This results in 6,343 test cases overall. For more details on the sampling process, see Appendix C. Conceptually, SimBenchGrouped measures the ability of LLMs to simulate responses from narrower participant groups based on specified group characteristics. [^4] Ideally, we would also like to measure LLM simulation ability for intersectional groups that combine multiple characteristics (e.g., female + age 30-49). However, selecting on multiple characteristics substantially decreases group size, thus increasing sampling noise in the response distributions. Reliable evaluation of intersectional group simulation ability would require datasets with more participants than we have access to.
@@ 3 Experimental Setup
Tested Models: To demonstrate the utility of SimBench and answer our six research questions (§1), we evaluate 45 recent LLMs. This includes both commercial and open-weight, base and instruction-tuned models, with sizes ranging from 0.5B to 405B parameters. Table 7 lists all models.
Model Elicitation: For each model, we collect predictions for the two main splits of SimBench2.4). To obtain model response distributions, we use one of two methods, depending on model type: 1) For base models, we directly extract token probabilities for each response option based on first-token logits. This is a natural way of eliciting a distribution out of an LLM, especially a base LLM. 2) For instruction-tuned models, we follow recent literature on LLM calibration and distribution prediction (Tian et al., 2023; Wang et al., 2024; Meister et al., 2025) and use verbalized distributions (e.g., “Option A: 30%, Option B: 70%”) elicited through prompting. We empirically validate this methodological choice in Appendix E, which provides strong evidence that verbalized distributions substantially and consistently outperform direct token probabilities for instruction-tuned models. This ensures each model class is evaluated using its most suitable answer elicitation method. For implementation details and prompt formats, see Appendix D.
Evaluation Metric: To measure LLM simulation ability, we derive the SimBench score $S$ from the Total Variation Distance (TVD). Conceptually, $S$ quantifies the improvement of a model’s prediction $Q$ over a uniform baseline $U$, relative to the human ground truth $P$:
$S(P,Q)=100\left(1-\frac{\operatorname{TVD}(P,Q)}{\operatorname{TVD}(P,U)}\right)$ (1)
where a score of 100 indicates perfect alignment and 0 indicates performance equivalent to random guessing. Since in practice a direct point-wise calculation is undefined when the question-level human distribution is uniform ($P=U$), to ensure numerical stability, we compute the score $S_{i}$ for each specific test case $i$ by normalizing against the dataset-level mean baseline:
$S_{i}=100\left(1-\frac{\operatorname{TVD}(P_{i},Q_{i})}{\frac{1}{|D|}\sum_{j\in D}\operatorname{TVD}(P_{j},U_{j})}\right)$ (2)
@@
where the denominator represents the average TVD between human responses and the uniform distribution across all test cases $j$ in the dataset $D$. This ensures the metric remains robust across datasets with varying entropy. The final reported score for a model is the average of $S_{i}$ across all evaluated test cases. As described in §2.4, we cap each dataset at 500 questions to limit over-representation of larger datasets; datasets with fewer total questions naturally contribute fewer test cases and thus carry proportionally less weight in the overall score.
4 Results
4.1 RQ1: General Simulation Ability of LLMs
To evaluate general simulation ability, we measure overall SimBench score $S$ averaged across the two main splits of SimBench in Table 1 and Appendix Table 7.
We find that the best LLMs achieve meaningful but modest simulation fidelity, as measured across the 20 datasets in SimBench. Claude-3.7-Sonnet is the best-performing model overall, achieving a score of 40.80 out of 100 on SimBench. While this score means the best model’s predictions remain closer to a uniform distribution than to the human ground truth, it nonetheless closes roughly 40% of the gap between the two, demonstrating a genuine and non-trivial simulation signal. The top open-weight model, DeepSeek-R1, scores 34.52. The majority of the 45 models we test perform substantially worse, scoring less than 20. Notably, ten models we test score below 0, indicating that their predicted response distributions are, on average, even further away from the true human response distribution than a uniform response distribution. Statistical analysis, detailed in Appendix H, confirms that the performance differences between most top-ranked models as well as within each model family are statistically significant. Overall, these results consolidate the mixed findings from prior work into a clearer, if somewhat sobering, picture. When evaluated across a diverse range of tasks and populations, today’s LLMs do possess genuine, non-trivial simulation abilities but are still far from being consistently reliable, general-purpose simulators. The stark performance differences between models also caution strongly against the use of smaller, less capable models for simulation, many of which perform worse than a simple uniform baseline.